Khalilsherman:

A newspaper reports the following changes in the temperature of a city over 4 days: 1.34 degrees Celsius, negative 5 over 7 degrees Celsius, −0.75 degrees Celsius, 4 over 9 degrees Celsius Write an expression to show the average daily change in the temperature of the city. Write the steps to solve the expression. Round your answer to the nearest hundredth\

Hoodmemes:

Ight so first u wanna see that the total of change in the temperature of the city Change of temp = 1.34 - 5/7 - 0.75 + 4/9

Hoodmemes:

And now u would do = 1.34 - 0.75 + (4/9 - 5/7)

Hoodmemes:

Next,=0.61 + (28/63 - 45/63),=0.61 - 17/63

Hoodmemes:

Last = 21.43/63

Hoodmemes:

Now what ya wanna do is divide \[(21.43/63)\div 4 = (blank)\]

Hoodmemes:

I-

Khalilsherman:

but imma try

Hoodmemes:

Ok um u have to be 13+ to be on this website u know that right?

Hoodmemes:

So what do u think the answer would be?

Khalilsherman:

0.85

Hoodmemes:

correct!

Hoodmemes:

I honestly would @ a moderator but im not a snitch and don condone snitching,so honestly i think u should delete your posts about your grade level plus age

Khalilsherman:

ok

Khalilsherman:

my fault im sorry

Hoodmemes:

its not,and its ok every body makes mistakes

Khalilsherman:

thank you for helping me

Hoodmemes:

np